Someday a person will walk on Mars, which is 5.6 x 1010 m from the earth at the closest approach. Determine the minimum time that will be required for a message from Mars to reach the earth via radio waves.

Question 1 options:


A) 

290 s


B) 

No option is correct


C) 

190 s


D) 

490 s


E) 

390 s



1
Expert's answer
2021-11-10T10:10:12-0500
"t=\dfrac{d}{c}=\dfrac{5.6\times10^{10}\ m}{3.0\times10^8\ \dfrac{m}{s}}=190\ s."

Therefore, the correct option is (C).

Answer:

C) 190 s